# Assumption-lean and Data-adaptive Post-Prediction Inference

### v3

@@ 5 Extensions
Here, we make extensions of POP-Inf to handle more general situations. In particular, Section 5.1 describes an inference procedure when the the outcome is fully labeled but covariate ${\bf x}$ is partly labeled and Section 5.2 considers the case where both $y$ and ${\bf x}$ are only partly labeled.
@@ 5.1 ML-predicted covariates
We adjusted the previously described setup by modifying it to use ML to predict the covariate ${\bf X}$ instead of the outcome $Y$. Our data can be divided into two parts: ${\cal L}\cup{\cal U}$. The ${\cal L}$ part includes data points $(y_{i},{\bf x}_{i},{\bf z}_{i})$ for $i=1,\cdots,n$ and the ${\cal U}$ part includes data points $(y_{i},{\bf z}_{i})$ for $i=n+1,\cdots,N+n$. In addition, we use an external prediction algorithm with the notation $\widehat{q}(\cdot)$. This algorithm is applied to ${\bf z}$ to generate the predicted value $\widehat{q}({\bf z})$ of the covariate ${\bf X}$. However, it is important to note that $\widehat{q}$ may produce inaccurate or biased predictions of ${\bf X}$. The estimand is also defined by the estimating equation (1).
Similar to our proposal to handle the ML-predicted outcome, we propose our POP-Inf estimator with ML-predicted covariates that solves the equation,
@@
$\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(y_{i},{\bf x}_{i};{\bm{\theta}})+\widehat{\bm{\omega}}\odot\left\{-\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(y_{i},\widehat{q}({\bf z}_{i});{\bm{\theta}})+\dfrac{1}{N}\sum_{i=n+1}^{n+N}\bm{\psi}(y_{i},\widehat{q}({\bf z}_{i});{\bm{\theta}})\right\}={\bf 0},$ (27)
where $\widehat{\bm{\omega}}$ is also to minimize the variance of the estimator. Such estimation equation is also the estimation equation for the classic estimator with an augmented mean zero term that is indexed by a vector $\bm{\omega}$. Therefore, the proposed estimator $\widehat{\bm{\theta}}_{\textnormal{POPInf}}(\bm{\omega})$ is also consistent and asymptotic normally distributed, under mild regularity conditions. Corollary 3 states the asymptotic distribution that enables interval estimation with its proof follows from the proof of Theorem 1.
@@
Under Conditions (C1)-(C3), assuming $\frac{n}{N}\rightarrow\rho$ as $n\to\infty$ and $N\rightarrow\infty$, then the proposed estimator $\widehat{\bm{\theta}}_{\textnormal{POPInf}}(\bm{\omega})$ converges to ${\bm{\theta}}_{0}$ in probability. Assuming additionally Condition (C4), we have that as $n\to\infty$ and $N\to\infty$,
@@
$\sqrt{n}(\widehat{\bm{\theta}}_{\textnormal{POP-Inf}}({\bm{\omega}})-{\bm{\theta}}_{0})\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}({\bf 0},{\bm{\Sigma}}^{\prime}(\bm{\omega})).$ (28)
@@
where ${\bm{\Sigma}}^{\prime}(\bm{\omega})={\bf A}^{-1}{\bf V}^{\prime}(\bm{\omega}){\bf A}^{-1}$, ${\bf V}^{\prime}(\bm{\omega})={\bf M}_{1}+\bm{\omega}\bm{\omega}^{\rm T}\odot({\bf M}_{2}^{\prime}+\rho{\bf M}_{3}^{\prime})-2{\bf I}_{q}\bm{\omega}^{\rm T}\odot{\bf M}_{4}^{\prime}$, ${\bf M}_{2}^{\prime}=\textnormal{Var}_{l}[\bm{\psi}(Y,\widehat{q}({\bf X});{\bm{\theta}}_{0})]$, ${\bf M}_{3}=\textnormal{Var}_{u}[\bm{\psi}(Y,\widehat{q}({\bf X});{\bm{\theta}}_{0})]$, ${\bf M}_{4}^{\prime}=\textnormal{Cov}_{l}[\bm{\psi}(Y,{\bf X};{\bm{\theta}}_{0}),\bm{\psi}(Y,\widehat{q}({\bf X});{\bm{\theta}}_{0})]$.

### v4

@@ 5 Extensions
Next, we extend PSPA to address more general scenarios. Specifically, Section 5.1 outlines an inference procedure for fully labeled outcomes with partially labeled covariate ${\bf x}$, and Section 5.2 discusses cases where both $y$ and ${\bf x}$ are partially labeled.
@@ 5.1 ML-predicted covariates
We adjusted the previously described setup by modifying it to use ML to predict the covariate ${\bf X}$ instead of the outcome $Y$. Our data can be divided into two parts: ${\cal L}\cup{\cal U}$. The ${\cal L}$ part includes data points $(y_{i},{\bf x}_{i},{\bf z}_{i})$ for $i=1,\cdots,n$ and the ${\cal U}$ part includes data points $(y_{i},{\bf z}_{i})$ for $i=n+1,\cdots,N+n$. In addition, we use an external prediction algorithm with the notation $\widehat{q}(\cdot)$. This algorithm is applied to ${\bf z}$ to generate the predicted value $\widehat{q}({\bf z})$ of the covariate ${\bf X}$. However, it is important to note that $\widehat{q}$ may produce inaccurate or biased predictions of ${\bf X}$. The estimand is also defined by the estimating equation (1). Similar to our proposal to handle the ML-predicted outcome, we propose our PSPA estimator $\widehat{\bm{\theta}}_{{\textnormal{{PSPA}}^{{}^{\prime}}}}({\bm{\omega}})$ with ML-predicted covariates that solves the equation
@@
$\Psi_{\textnormal{{PSPA}}*}^{\bm{\omega}}({\bm{\theta}}):=\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(y_{i},{\bf x}_{i};{\bm{\theta}})+{\rm diag}({\bm{\omega}})\left\{-\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(y_{i},\widehat{q};{\bm{\theta}})+\dfrac{1}{N}\sum_{i=n+1}^{n+N}\bm{\psi}(y_{i},\widehat{q};{\bm{\theta}})\right\}={\bf 0},$ (26)
Denote ${\bf A}=\mathbb{E}[\partial\bm{\psi}(Y,{\bf X};{\bm{\theta}})/\partial{\bm{\theta}}]$, ${\bf M}_{1}=\operatorname{Var}[\bm{\psi}(Y,{\bf X};{\bm{\theta}})]$, ${\bf M}_{2}^{\prime}=\operatorname{Var}[\bm{\psi}(Y,\widehat{q};{\bm{\theta}}_{0})]$, ${\bf M}_{3}^{\prime}=\operatorname{Var}[\bm{\psi}(Y,\widehat{q};{\bm{\theta}}_{0})]$, ${\bf M}_{4}^{\prime}=\operatorname{Cov}[\bm{\psi}(Y,{\bf X};{\bm{\theta}}_{0}),\bm{\psi}(Y,\widehat{q};{\bm{\theta}}_{0})]$. Upon calculations similar to Theorem 1, the $j$-th diagonal element of the asymptotic variance of $\widehat{\bm{\theta}}_{{\textnormal{{PSPA}}^{{}^{\prime}}}}({\bm{\omega}})$ is
${\bm{\Sigma}}^{\prime}_{jj}({\bm{\omega}})$ $=\omega_{j}^{2}[{\bf A}^{-1}({\bf M}_{2}^{\prime}+\rho{\bf M}_{3}^{\prime}){\bf A}^{-1}]_{jj}-2\omega_{j}[{\bf A}^{-1}{\bf M}_{4}^{\prime}{\bf A}^{-1}]_{jj}+[{\bf A}^{-1}{\bf M}_{1}^{\prime}{\bf A}^{-1}]_{jj},$ (27)
which achieve its minimum with $\omega_{j}^{\textnormal{opt}}=\frac{[{\bf A}^{-1}{\bf M}_{4}^{\prime}{\bf A}^{-1}]_{jj}}{[{\bf A}^{-1}({\bf M}_{2}^{\prime}+\rho{\bf M}_{3}^{\prime}){\bf A}^{-1}]_{jj}}$. We modify our Algorithm 1 to the following algorithm to incorporate ML-predicted covariates.
Algorithm 2 PSPA estimation with ML-predicted covariates
1:Data ${\cal L}\cup{\cal U}$, pre-trained ML model $\widehat{f}$, error rate $\alpha\in(0,1)$.
2:Obtain the classical estimator $\widehat{\bm{\theta}}_{\text{C}}$ by solving $\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(y_{i},{\bf x}_{i};{\bm{\theta}})={\bf 0}$.
3:Obtain the optimal weighting vector $\widehat{\bm{\omega}}^{\textnormal{opt}}=[\widehat{\omega}_{1}^{\textnormal{opt}},\dots,\widehat{\omega}_{q}^{\textnormal{opt}}]^{\rm T}$ by
$\widehat{\omega}_{j}^{\textnormal{opt}}=\begin{cases}\min(\frac{[\widehat{\bf A}_{\text{C}}^{-1}\widehat{\bf M}_{4,{\text{C}}}^{{}^{\prime}}\widehat{\bf A}_{\text{C}}^{-1}]_{jj}}{[\widehat{\bf A}_{\text{C}}^{-1}(\widehat{\bf M}_{2,{\text{C}}}^{{}^{\prime}}+\rho\widehat{\bf M}_{3,{\text{C}}}^{{}^{\prime}})\widehat{\bf A}_{\text{C}}^{-1}]_{jj}},\frac{\lambda_{{\min,+}}[\frac{1}{n}\nabla\bm{\psi}(y_{i},{\bf x}_{i};{\bm{\theta}})]}{\lambda_{\max}[\frac{1}{n}\nabla\bm{\psi}(y_{i},\widehat{q};{\bm{\theta}})]})&\quad\text{if }\frac{[\widehat{\bf A}_{\text{C}}^{-1}\widehat{\bf M}_{4,{\text{C}}}^{{}^{\prime}}\widehat{\bf A}^{-1}]_{jj}}{[\widehat{\bf A}_{\text{C}}^{-1}(\widehat{\bf M}_{2,{\text{C}}}^{{}^{\prime}}+\rho\widehat{\bf M}_{3,{\text{C}}}^{{}^{\prime}})\widehat{\bf A}_{\text{C}}^{-1}]_{jj}}>0\\ \max(\frac{[\widehat{\bf A}_{\text{C}}^{-1}\widehat{\bf M}_{4,{\text{C}}}^{{}^{\prime}}\widehat{\bf A}_{\text{C}}^{-1}]_{jj}}{[\widehat{\bf A}_{\text{C}}^{-1}(\widehat{\bf M}_{2,{\text{C}}}^{{}^{\prime}}+\rho\widehat{\bf M}_{3,{\text{C}}}^{{}^{\prime}})\widehat{\bf A}_{\text{C}}^{-1}]_{jj}},0)&\quad\text{if }\frac{[\widehat{\bf A}_{\text{C}}^{-1}\widehat{\bf M}_{4,{\text{C}}}^{{}^{\prime}}\widehat{\bf A}^{-1}]_{jj}}{[\widehat{\bf A}_{\text{C}}^{-1}(\widehat{\bf M}_{2,{\text{C}}}^{{}^{\prime}}+\rho\widehat{\bf M}_{3,{\text{C}}}^{{}^{\prime}})\widehat{\bf A}_{\text{C}}^{-1}]_{jj}}\leq 0\\ \end{cases}$
for all $j\in 1,\dots,q$. Here, $\widehat{\bf A}_{\text{C}}$, $\widehat{\bf M}_{1,{\text{C}}}^{{}^{\prime}}$, $\widehat{\bf M}_{2,{\text{C}}}^{{}^{\prime}}$, $\widehat{\bf M}_{3,{\text{C}}}^{{}^{\prime}}$, $\widehat{\bf M}_{4,{\text{C}}}^{{}^{\prime}}$ are sample analogs of ${\bf A}$, ${\bf M}_{1}^{{}^{\prime}}$, ${\bf M}_{2}^{{}^{\prime}}$, ${\bf M}_{3}^{{}^{\prime}}$, ${\bf M}_{4}^{{}^{\prime}}$ for $\Psi_{\textnormal{{PSPA}}*}^{\bm{\omega}}({\bm{\theta}})$ with $\widehat{\bm{\theta}}_{\text{C}}$ plugged in. $\lambda_{{\min,+}}[{\bf M}]$ and $\lambda_{\max}[{\bf M}]$ represent the smallest non-negative and largest eigenvalue of matrix ${\bf M}$, respectively.
4:Obtain the $\textnormal{{PSPA}}^{*}$ estimator with a one-step update:
$\widehat{\bm{\theta}}_{{\textnormal{{PSPA}}^{{}^{\prime}}}}=\widehat{\bm{\theta}}_{\text{C}}-\left[\nabla\Psi_{\textnormal{{PSPA}}^{{}^{\prime}}}^{\widehat{\bm{\omega}}^{\textnormal{opt}}}(\widehat{\bm{\theta}}_{\text{C}})\right]^{-1}\Psi_{\textnormal{{PSPA}}^{{}^{\prime}}}^{\widehat{\bm{\omega}}^{\textnormal{opt}}}(\widehat{\bm{\theta}}_{\text{C}}),\text{ where }\nabla\Psi_{\textnormal{{PSPA}}^{{}^{\prime}}}^{\widehat{\bm{\omega}}^{\textnormal{opt}}}=\partial\Psi_{\textnormal{{PSPA}}^{{}^{\prime}}}^{\widehat{\bm{\omega}}^{\textnormal{opt}}}({\bm{\theta}})/\partial{\bm{\theta}}|_{{\bm{\theta}}=\widehat{\bm{\theta}}_{\text{C}}}.$
5:Obtain the asymptotic variance of $\widehat{\bm{\theta}}_{{\textnormal{{PSPA}}^{{}^{\prime}}}}$:
$\widehat{\bm{\Sigma}}^{\prime}(\widehat{\bm{\omega}}^{\textnormal{opt}})=\widehat{\bf A}_{\textnormal{{PSPA}}^{{}^{\prime}}}^{-1}[\widehat{\bf M}_{1,\textnormal{{PSPA}}}^{{}^{\prime}}+{\rm diag}(\widehat{\bm{\omega}}^{\textnormal{opt}})(\widehat{\bf M}_{2,\textnormal{{PSPA}}}^{{}^{\prime}}+\rho\widehat{\bf M}_{3,\textnormal{{PSPA}}}^{{}^{\prime}}){\rm diag}({\bm{\omega}})-2{\rm diag}(\widehat{\bm{\omega}}^{\textnormal{opt}})\widehat{\bf M}_{4,\textnormal{{PSPA}}}^{{}^{\prime}}]\widehat{\bf A}_{\textnormal{{PSPA}}^{{}^{\prime}}}^{-1},$
where $\widehat{\bf A}_{\textnormal{{PSPA}}}^{{}^{\prime}}$, $\widehat{\bf M}_{1,\textnormal{{PSPA}}}^{{}^{\prime}}$, $\widehat{\bf M}_{2,\textnormal{{PSPA}}}^{{}^{\prime}}$, $\widehat{\bf M}_{3,\textnormal{{PSPA}}}^{{}^{\prime}}$, $\widehat{\bf M}_{4,\textnormal{{PSPA}}}^{{}^{\prime}}$ are sample analogs of ${\bf A}$, ${\bf M}_{1}^{{}^{\prime}}$, ${\bf M}_{2}^{{}^{\prime}}$, ${\bf M}_{3}^{{}^{\prime}}$, ${\bf M}_{4}^{{}^{\prime}}$ for $\Psi_{\textnormal{{PSPA}}^{{}^{\prime}}}^{\bm{\omega}}({\bm{\theta}})$ with $\widehat{\bm{\theta}}_{\textnormal{{PSPA}}}$ plugged in.
6:$\textnormal{{PSPA}}^{{}^{\prime}}$ estimator $\widehat{\bm{\theta}}_{{\textnormal{{PSPA}}^{{}^{\prime}}}}$, standard error $\sqrt{\frac{\widehat{\bm{\Sigma}}^{\prime}(\widehat{\bm{\omega}}^{\textnormal{opt}})_{jj}}{n}}$, $\alpha$-level confidence interval $\mathcal{C}_{\alpha,j}^{\textnormal{{PSPA}}^{{}^{\prime}}}=(\widehat{\bm{\theta}}_{\textnormal{{PSPA}}^{{}^{\prime}}_{j}}\pm z_{1-\alpha/2}\sqrt{\frac{\widehat{\bm{\Sigma}}^{\prime}(\widehat{\bm{\omega}}^{\textnormal{opt}})_{jj}}{n}})$, and (two-sided) p-value $2(1-\Phi(|\frac{\widehat{{\bm{\theta}}}_{\textnormal{{PSPA}}^{{}^{\prime}}_{j}}}{\sqrt{\frac{\widehat{\bm{\Sigma}}^{\prime}(\widehat{\bm{\omega}}^{\textnormal{opt}})_{jj}}{n}})}|))$ for the $j$-th coordinate. Here, $\Phi$ is the CDF of the standard normal distribution.
Corollary 4 states the asymptotic distribution that enables interval estimation with its proof follows from the proof of Theorem 1, Corollary 1 and 2, and thus omitted.
@@
Under Conditions (C1), (C2) on equation 26, and (C3)-(C4), assuming $\frac{n}{N}\rightarrow\rho$ as $n\to\infty$ and $N\rightarrow\infty$, and $\widehat{\bm{\omega}}\stackrel{{\scriptstyle P}}{{\rightarrow}}{\bm{\omega}}$, then the proposed estimator $\widehat{\bm{\theta}}_{\textnormal{{PSPA}}}(\widehat{\bm{\omega}})$ converges to ${\bm{\theta}}_{0}$ in probability. Assuming additionally Condition (C5), we have that as $n\to\infty$ and $N\to\infty$,
@@
$\sqrt{n}(\widehat{\bm{\theta}}^{\prime}_{\textnormal{{{PSPA}}}}({\widehat{\bm{\omega}}})-{\bm{\theta}}_{0})\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}({\bf 0},{\bm{\Sigma}}^{\prime}({\bm{\omega}})).$ (28)
@@
where ${\bm{\Sigma}}^{\prime}({\bm{\omega}})={\bf A}^{-1}{\bf V}^{\prime}({\bm{\omega}}){\bf A}^{-1}$, ${\bf A}=\mathbb{E}[\partial\bm{\psi}(Y,{\bf X};{\bm{\theta}})/\partial{\bm{\theta}}]$, ${\bf V}^{\prime}({\bm{\omega}})={\bf M}_{1}+{\rm diag}({\bm{\omega}})({\bf M}_{2}^{\prime}+\rho{\bf M}_{3}^{\prime}){\rm diag}({\bm{\omega}})-2{\rm diag}({\bm{\omega}}){\bf M}_{4}^{\prime}$, ${\bf M}_{1}=\operatorname{Var}[\bm{\psi}(Y,{\bf X};{\bm{\theta}})]$, ${\bf M}_{2}^{\prime}=\operatorname{Var}[\bm{\psi}(Y,\widehat{q};{\bm{\theta}}_{0})]$, ${\bf M}_{3}^{\prime}=\operatorname{Var}[\bm{\psi}(Y,\widehat{q};{\bm{\theta}}_{0})]$,
${\bf M}_{4}^{\prime}=\operatorname{Cov}[\bm{\psi}(Y,{\bf X};{\bm{\theta}}_{0}),\bm{\psi}(Y,\widehat{q};{\bm{\theta}}_{0})]$. With $\widehat{\bm{\Sigma}}^{\prime}(\widehat{\bm{\omega}})\stackrel{{\scriptstyle P}}{{\rightarrow}}{\bm{\Sigma}}^{\prime}({\bm{\omega}})$,
$\lim_{n\rightarrow\infty}\mathbb{P}(\theta_{j}\in\mathcal{C}_{\alpha,j}^{{{}^{\prime}}\widehat{\bm{\omega}}})\geq 1-\alpha,\ j=1,\cdots,q,$ (29)
where $\mathcal{C}_{\alpha,j}^{{{}^{\prime}}\widehat{\bm{\omega}}}=(\widehat{\bm{\theta}}_{\textnormal{{{PSPA}}}_{j}}^{\prime}(\widehat{\bm{\omega}})\pm z_{1-\alpha/2}\sqrt{\frac{\widehat{\bm{\Sigma}}^{\prime}(\widehat{\bm{\omega}})_{jj}}{n}})$, $1-\alpha\in(0,1)$ is the confidence level, and $\theta_{j}$ is the $j$-th coordinate of the parameter.
Remark 3.
Condition (C2) requires equation 26 to have a unique solution. To ensure this, we impose a constraint on ${\bm{\omega}}$. Equation 26 has a unique solution if its derivative, $\dfrac{1}{n}\sum_{i=1}^{n}[\nabla\bm{\psi}(y_{i},{\bf x}_{i};{\bm{\theta}})-{\rm diag}({\bm{\omega}})\nabla\bm{\psi}(y_{i},\widehat{q};{\bm{\theta}})]+\dfrac{1}{N}\sum_{i=n+1}^{n+N}\nabla\bm{\psi}(y_{i},\widehat{q};{\bm{\theta}}),$ is positive semi-definite. The second term, $\dfrac{1}{N}\sum_{i=n+1}^{n+N}\nabla\bm{\psi}(y_{i},\widehat{q};{\bm{\theta}})$, is already positive semi-definite, so we only need to ensure the first term is positive semi-definite. According to Lemma 2 in the appendix, a sufficient condition for this is: $0\leq\omega_{j}\leq\dfrac{\lambda_{{\min,+}}[\dfrac{1}{n}\sum_{i=1}^{n}\nabla\bm{\psi}(y_{i},{\bf x}_{i};{\bm{\theta}})]}{\lambda_{\max}[\dfrac{1}{n}\sum_{i=1}^{n}\nabla\bm{\psi}(y_{i},\widehat{q};{\bm{\theta}})]}$ for all $j=1,\cdots,q$, where $\lambda_{{\min,+}}[{\bf M}]$ and $\lambda_{\max}[{\bf M}]$ represent the smallest positive and largest eigenvalue of matrix ${\bf M}$, respectively. By imposing this constraint on the weighting vector in the algorithm, we ensure that condition (C2) holds. In practice, one can relax this constraint by substituting the prefixed ${\bm{\omega}}$ into the derivative and confirming its positive semi-definiteness. If necessary, one can iteratively reduce the value of each positive $\omega_{j}$ (or increase negative $\omega_{j}$)w until the derivative achieves positive semi-definiteness.
Next, we state the theoretical guarantees of PSPA for element-wise variance reduction compared with the classical approach with weighting vector ${\bm{\omega}}^{\textnormal{C}}=[0,\dots,0]^{\rm T}$.
Corollary 5.
Suppose $\sqrt{n}(\widehat{\bm{\theta}}_{\textnormal{{PSPA}}^{\prime}}({\bm{\omega}}^{\textnormal{C}})-{\bm{\theta}})\stackrel{{\scriptstyle D}}{{\rightarrow}}N({\bf 0},{\bm{\Sigma}}({\bf 0}))$. Denote $\hat{\bm{\omega}}^{\textnormal{opt}}=[\widehat{\omega}_{1}^{\textnormal{opt}},\dots,\widehat{\omega}_{q}^{\textnormal{opt}}]^{\rm T}$, where
$\widehat{\omega}_{j}^{\textnormal{opt}}=\begin{cases}\min(\frac{[\widehat{\bf A}^{-1}\widehat{\bf M}_{4,{\text{C}}}^{{}^{\prime}}\widehat{\bf A}^{-1}]_{jj}}{[\widehat{\bf A}^{-1}(\widehat{\bf M}_{2,{\text{C}}}^{{}^{\prime}}+\rho\widehat{\bf M}_{3,{\text{C}}}^{{}^{\prime}})\widehat{\bf A}^{-1}]_{jj}},\frac{\lambda_{{\min,+}}[\frac{1}{n}\nabla\bm{\psi}(y_{i},{\bf x}_{i};{\bm{\theta}})]}{\lambda_{\max}[\frac{1}{n}\nabla\bm{\psi}(y_{i},\widehat{q};{\bm{\theta}})]})&\quad\text{if }\frac{[\widehat{\bf A}^{-1}\widehat{\bf M}_{4,{\text{C}}}^{{}^{\prime}}\widehat{\bf A}^{-1}]_{jj}}{[\widehat{\bf A}^{-1}(\widehat{\bf M}_{2,{\text{C}}}^{{}^{\prime}}+\rho\widehat{\bf M}_{3,{\text{C}}}^{{}^{\prime}})\widehat{\bf A}^{-1}]_{jj}}>0\\ \max(\frac{[\widehat{\bf A}^{-1}\widehat{\bf M}_{4,{\text{C}}}^{{}^{\prime}}\widehat{\bf A}^{-1}]_{jj}}{[\widehat{\bf A}^{-1}(\widehat{\bf M}_{2,{\text{C}}}^{{}^{\prime}}+\rho\widehat{\bf M}_{3,{\text{C}}}^{{}^{\prime}})\widehat{\bf A}^{-1}]_{jj}},0)&\quad\text{if }\frac{[\widehat{\bf A}^{-1}\widehat{\bf M}_{4,{\text{C}}}^{{}^{\prime}}\widehat{\bf A}^{-1}]_{jj}}{[\widehat{\bf A}^{-1}(\widehat{\bf M}_{2,{\text{C}}}^{{}^{\prime}}+\rho\widehat{\bf M}_{3,{\text{C}}}^{{}^{\prime}})\widehat{\bf A}^{-1}]_{jj}}\leq 0\\ \end{cases}$ (30)
for all $j\in 1,\dots,q$. Suppose $\hat{\bm{\omega}}^{\textnormal{opt}}\stackrel{{\scriptstyle P}}{{\rightarrow}}{\bm{\omega}}^{\textnormal{opt}}$, then ${\bm{\Sigma}}({\bm{\omega}}^{\textnormal{opt}})_{jj}\leq{\bm{\Sigma}}({\bm{\omega}}^{\textnormal{C}})_{jj}$.
The proof is contained in Appendix B.6.